OpenStudy (anonymous):

Multiply and simplify. x-1/x^2+2x+1 multiplied by x^2-1/x+1 1 correct? (x-1)^2/(x+1)^2 x-1/x+1 x+1/x-1

hero (hero):

Why do you believe 1 is correct? Do you mind showing your work?

OpenStudy (anonymous):

I crossed multiplied and it cancelled out

hero (hero):

You can only cross multiply when you have something like \(\dfrac{a}{b} = \dfrac{c}{d}\)

OpenStudy (anonymous):

Is my answer correct?

hero (hero):

You have to factor and cancel to simplify

OpenStudy (anonymous):

so my answer is wrong. How to factor and simplify?

hero (hero):

Write the trinomial in the form (x + a)(x + b) Also \(x^2 - 1\) is a difference of squares. Remember \(a^2 - b^2 = (a + b)(a - b)\)

OpenStudy (anonymous):

I'm not good at this, x2-1 + x-1?